Please use Rstudio to answer the following question. Thank you.

In a double-blind experiment, healthy volunteers agreed to be exposed to a cold virus. They were either assigned to a control group (no echinacea but a placebo) or the group with echinacea. Here are the results Placebo group: 103 subjects 88 got a cold Echinacea group: 48 subjects 44 got a cold.

a. Calculate the proportions getting cold in each group.

b. Use a two-sample test to compare the proportions. In this case, use a continuity correction.

c. Give a 95% confidence interval for the difference comparing echinacea and placebo.

d. Comment on possible odditites in the study design?  
Answer #1

a)

Given,

Placebo group: n1 103, r1-88

Sample proportion, Pi = n1 = 803 = 0.8544 n 103

Echinacea group: n2= 48, x2= 44

Sample proportion, P2 = n2 = 48 48-0.9167 = 0.9167 P2 =

-------------------------------------------------------------------------------------------------------------

b)

Our hypothesis is,

Ho:P-P2

H_1: P_1 eq P_2

Test statistics:

P1-P2 RI

Where, p1288+44 n1+no 103+48 = 0.8742

P1-P2 0.8544-0.9167 POG+ 0.0623 ) 、/0.8742 * 0.1258(T- 010348 T 0.05795554 | 1.074962| 1.075

Using continuity correction subtract 0.5 from the absolute value i.e. 1.075-0.5 = 0.575

At 0.05 level of significance, critical z value is 1.96

Since, calculated z value is less than the critical z value, i.e.,

calcrulated1.96, Ho is not rejected s not reecte

Therefore, we conclude that the proportion of two groups are insignificant.

-------------------------------------------------------------------------------------------------------------------------------

c).

95% confidence interval for Pi - Pz

n1 n2

0.0623±1.96*0.05795554ーー0.0623±0.1135929= (-0.1758929. 0.0512929) つつつ